OpenStudy (anonymous):

The equation x2 + y2 − 2x + 2y − 1 = 0 is the general form of the equation of a circle. What is the standard form of the equation?

OpenStudy (anonymous):

(x − 1)2 − (y − 1)2 = 1 (x − 1)2 + (y + 1)2 = 3 (x + 1)2 + (y − 1)2 = 1 (x − 1)2 + (y + 1)2 = 1

OpenStudy (phi):

Do you know how to "complete the square" ? See https://www.khanacademy.org/math/algebra/quadratics/completing_the_square/v/completing-the-square for how to do it. I would complete the square on x^2 − 2x and complete the square on y^2 +2y